Parth (parthkohli):

Prove:\[1 < \frac{1}{1001} + \frac{1}{1002} + \frac{1}{1003} + \cdots + \frac{1}{3001} < 1 \frac{1}3\]

Parth (parthkohli):

The left inequality is a standard AM-HM one. How about the right?

Parth (parthkohli):

@ganeshie8

Parth (parthkohli):

Calc seems to work. Algebra?

OpenStudy (mathmate):

Since the required sum is only about 1.0986 with a big margin from 4/3, we can brute force the right inequality: Sum < 500(1/1001+1/1501+1/2001+1/2501)=1.3263<4/3

OpenStudy (mathmate):

Here's a more refined proof for the second inequality: We define \(s(m)=\frac{1}{1001+m}+\frac{1}{3001-m} \) and it can be shown that \(s(i)>s(i+1) \) for \(0\le i\le 999\).........(1) \(S=\sum_{k=1001}^{k=3001}1/k\) \(=(\sum_{m=0}^{m=1000}s(m) ~) ~~-1/2001\) \(< 1001s(0)-1/2001 \)...............using (1) \(= 1.33305573 \) \(< 4/3\) Proof for (1): \(s(n)-s(n+1))=-1/(n+1002)+1/(n+1001)+1/(3001-n)-1/(3000-n)\) \(=1/(n^2+2003n+1003002)-1/(n^2-6001n+9003000)\) \(>0~\forall ~0\le n\le 999\)